OpenStudy (anonymous):

1. y=x+6 y=2x a.(6,12) b.(-12,-6) c.(-6,-12) d.(2,4)

OpenStudy (anonymous):

You need to find the set of points that makes both equations true. A is the only one that is true for both. If you need a little more explanation I'd be happy to help!

OpenStudy (anonymous):

can you help me on one more

OpenStudy (anonymous):

Sure thing! Post it and I'll do my best, since I wouldn't call myself a math genius or anything ^^

OpenStudy (anonymous):

2. 3x+2y=7 -3x+11+4 a.(6,-3) b.(6,7) c.(-4,19/2) d.(5.-4)

OpenStudy (anonymous):

You would use the same strategy and plug in the points into both equations. Whichever set of points works for both equations is the correct answer. Only issue here is that I don't understand the second equation. Are you sure you wrote it correctly?

OpenStudy (anonymous):

to find x set y to 0. then solve

OpenStudy (anonymous):

3x+2y=7 -3x+11=y a.(6,-3) b.(6,7) c.(-4,19/2) d.(5,-4)

OpenStudy (anonymous):

Okay that looks better. Just do the same as the first problem and plug in the various points to see which one works for both equations. You should get D to be the correct answer.

OpenStudy (anonymous):

thx so much